# Specialization after Generalization: Towards Understanding Test-Time Training in Foundation Models

### v1

@@ 1 Introduction
We propose to model this phenomenon under the linear representation hypothesis (LRH; Mikolov et al., 2013; Park et al., 2024; 2025), which postulates that models represent high-level concepts—meaningful semantic features—as directions in a latent space. This latent concept space is typically assumed to be sparse conditioned on the input, meaning that any given input activates only a few possible concepts. For a particular prediction task, the linear coefficients of the active concepts effectively characterize their “meaning” for that task. Because the number of real-world concepts far exceeds a model’s dimensionality, these concepts are superimposed within the model’s dense activations (cf. Figure 1; Elhage et al., 2022). The LRH has been used extensively in prior work on interpretability (Kim et al., 2018) and activation steering (Bolukbasi et al., 2016; Templeton et al., 2024) of foundation models. In this work, we analyze a model where TTT can learn the meaning of these superimposed concepts from data more efficiently than training a “global” model or non-parametric methods.
@@
The LRH can be formalized as follows. We assume the existence of an $s$-sparse concept space $\Phi:\mathcal{X}\to\mathbb{R}^{d_{1}}$} which is approximated by a learned feature map $\Psi:\mathcal{X}\to\mathbb{R}^{d_{2}}$} with $d_{2}\ll d_{1}$. The ground truth to be learned is linear in the concept space, i.e., $f(x)=\langle\Phi(x),w_{\star}\rangle$ for some unknown $w_{\star}\in\mathbb{R}^{d_{1}}$. [^1] In classification, the logits are canonically parameterized as $\langle\Phi(x),w_{c}\rangle$ where $w_{c}$ is the class-specific weight vector. The class probability is then given by $\mathbb{P}(c\mid x)\propto\exp(\langle\Phi(x),w_{c}\rangle)$. In Section 3, we leverage the LRH to develop a mechanistic understanding of how TTT behaves, and make the following key observations:
@@ 3. O3:
Based on the LRH and our key observations, in Sections 4 and 5, we analyze when and why TTT can be effective. We find that TTT improves predictions in the underparameterized regime, but its improvement diminishes as models become overparameterized. We support this finding empirically and theoretically:
@@ 2 Related work
So far it has not been well understood why and when TTT is effective. While many different methods have been proposed for TTT, we focus here on analyzing “semi-parametric” TTT (e.g., Hardt & Sun, 2024; Hübotter et al., 2025), where a pre-trained model is fine-tuned with a supervised loss on a small neighborhood of the test point in the training data. This is different from some other methods for test-time “adaptation”, which are commonly applied with distribution shifts (e.g., Wang et al., 2021; Zhang et al., 2022; Durasov et al., 2025). Basu et al. (2023) consider a similar setting to ours, but analyze it through the lens of non-parametric estimation, relying on the smoothness of the target function in the feature space $\Psi$. In contrast, our framework explicitly models the underlying sparse concept space $\Phi$. This explains why TTT substantially outperforms “non-parametric” methods even when the function is locally high-dimensional ($s$-sparse) in the concept space. Furthermore, while most prior theoretical work simply assumes the TTT gradient aligns with the gradient on the oracle label (e.g., Sun et al., 2020), our work provides an idealized model where this alignment is justified.
@@ 3 How does specialization behave?
We begin in this section by developing a mechanistic understanding of how TTT behaves. Since we do not have access to the “true” hypothesized concept space $\Phi$, we train SAEs to learn a concept space $\smash{\hat{\Phi}}$ whose characteristics can be probed. We employ a top-$k$ SAE (Gao et al., 2025) to learn sparse feature representations. After a brief description of the experimental setup, we detail our key observations O1-O3 from Section 1.
@@ 3.1 Experimental setup
We use the ImageNet-1K dataset (Deng et al., 2009). The dense vectors $\Psi(x)$ are normalized CLIP embeddings (Radford et al., 2021) of dimension $d_{2}=512$ (we use only the <CLS> component). We set the sparse dimension to $d_{1}=8\times d_{2}=4096$ and the sparsity level to $s=16$. For our analysis, we use the SAE’s reconstructions $\smash{\hat{\Psi}(x)}$ rather than the original CLIP embeddings $\Psi(x)$. This choice aligns our experiments more closely with our theoretical model and circumvents known challenges in training SAEs on raw, complex embeddings. This comes at the cost of a mild $6\%$ drop in accuracy for a global linear classifier trained on the embeddings. In Section E.2, we additionally validate the SAE framework on MNIST (LeCun et al., 1998b), where we use $\Psi(x)$ directly rather than its reconstruction, as recovering an accurate $\smash{\hat{\Phi}(x)}$ is much easier on this simpler dataset. We include training details in Section E.3.
@@ 3.2 Results
Our first hypothesis is that the SAE mapping preserves the angular relationships between a point and its neighbors. To test this, we select a neighborhood for a test point $x^{\star}$ in three different spaces: the original CLIP space ($\Psi$), the reconstructed space ($\smash{\hat{\Psi}}$), and the estimated concept space ($\smash{\hat{\Phi}}$). We then measure the average cosine similarity in the estimated concept space between $x^{\star}$ and points in each neighborhood. As shown in Figure 2, the distributions of cosine similarities are nearly identical regardless of the space used for neighbor selection. This suggests that the SAE projection to the concept space preserves the local geometric structure, supporting our first key observation.
@@
O2: Neighborhoods are supported by few concepts.
With a sparsity penalty of $\lambda=0.2$, the learned masks are highly sparse, activating on average only $\|m\|_{0}\approx 40$ concepts. This is substantially smaller than the total number of unique concepts active across the neighborhood, which is approximately $180$. As shown in Table 1 (TTT column), this sparsely supported model performs on par with TTT on top of dense reconstructions $\smash{\hat{\Psi}(x)}$ [^4] TTT in the concept space without extra masking achieves similar results. The fact that it also works with masking provides strong evidence that TTT adjusts only a few selected concepts in the neighborhood.. This suggests that a small, adaptively chosen set of concepts is sufficient to capture the relevant information within a local region. We obtain similar results for the Gemma Scope SAE (Lieberum et al., 2024) on MNIST data, which we present in Section E.2.
@@
Figure 3: Comparison of predicted class probabilities for TTT models trained on dense reconstructions ($\smash{\hat{\Psi}}$) and sparse concepts ($\smash{\hat{\Phi}}$), relative to their magnitude. The small relative TV distance ($\ll\!\!1$; defined in Section E.1) between both distributions indicates strong functional agreement between TTT in $\smash{\hat{\Psi}}$ and $\smash{\hat{\Phi}}$ space. We show 90% bootstrap confidence intervals across $1000$ test points.
@@
While the adaptive masking in Equation 3 explicitly enforces a sparse solution, we find evidence that standard TTT in the feature space implicitly favors a solution that is sparse in the concept space. First, the TTT models trained on dense reconstructions $\smash{\hat{\Psi}(x)}$ and sparse concepts $\smash{\hat{\Phi}_{m}(x)}$ achieve nearly identical accuracy (cf. Table 1). Furthermore, their predictions agree in $\approx\!\!89\%$ of cases, indicating that they learn functionally equivalent classifiers (apart from pathological examples). Figure 3 reinforces this by showing that both models lead to closely matched predictive distributions over the top-10 predicted classes. In Figure 3, we compare the ordered predicted probabilities for $\smash{\hat{\Phi}}$ to the corresponding probabilities for $\smash{\hat{\Psi}}$, matching the distributions’ temperatures, and averaging over all test points. Their strong correspondence suggests that TTT on reconstructed embeddings is implicitly biased towards a sparse solution in the underlying concept space. This phenomenon may be linked to the implicit bias of optimization algorithms (e.g., SGD or Adam), which are known to favor minimum-norm solutions (Gunasekar et al., 2018). When the feature map superimposes concepts, this implicit bias may favor sparse solutions in the underlying concept space (Vaskevicius et al., 2019).
@@ 4 When does specialization help?
After gaining some mechanistic understanding of TTT in Section 3, we next study when specialization through TTT improves over a globally trained model. In Section 5, we then relate our results from §3 and §4, by providing theoretical support for why the LRH and our observations O1-O3 may support our practically relevant findings from this section.
@@ 4.1 Settings
We consider the following three tasks, and provide implementation details in Appendix E:
  • MNIST. We train a global classifier following the LeNet-5 (LeCun et al., 1998a) architecture with a cross-entropy loss on the MNIST (LeCun et al., 1998b) dataset of handwritten digits. To distinguish learning of concepts from TTT’s ability to identify the correct weights for each locally relevant concept, we train only the last linear layer during TTT. The last linear layer is updated to minimize the cross-entropy loss on the neighborhood of a test point (cf. Equation 2). We report classification error on the test set.
@@ 4.2 Results
We perform a scaling study by varying the model size and comparing the performance of TTT against global training and a majority vote on the neighborhood (i.e., a simple non-parametric baseline). Figure 4 shows the results. In MNIST, we train convolutional neural networks of varying sizes as summarized in Appendix E. In ImageNet, we train multi-layer perceptrons on top of CLIP embeddings, varying the hidden dimension. In language modeling, we evaluate Qwen2.5 base models of sizes ranging from 0.5B to 32B parameters. We find across all tasks that TTT outperforms global training and majority vote, with the performance gap shrinking as the model size increases. We hypothesize that at a larger model size, fewer concepts have to be superimposed in latent space, leading to less interference when globally mapping latent representations to predictions. While a larger model size allows for better global disentanglement of concepts, TTT can compensate for limited model capacity by adapting the head to the specific concepts in a local neighborhood.
@@
Next to performing a scaling study on model size, we also vary the dataset size. Figure 5 shows the results at a fixed model scale. We subsample the training datasets of ImageNet and MNIST to fractions ranging from $1\%$ to $100\%$ of the original training set size, ensuring that all subsampled datasets are class-balanced. We then train global models and evaluate TTT on the respective test sets. We find that TTT consistently outperforms global training, with the slight trend of the performance gap widening as the dataset size increases. We hypothesize that larger datasets provide richer neighborhoods for local adaptation, enabling TTT to specialize more effectively to the specific concepts relevant to each test point.
@@
To validate that TTT improves predictions locally, we globally evaluate some TTT heads. As shown in Table 2, the global accuracy of these fixed TTT heads is significantly lower than that of the model trained on the entire dataset without TTT. These results confirm that, while TTT can provide localized performance improvements when adapted individually at test time, such benefits do not generalize when the same adaptation is applied globally. We further hypothesize that the neighborhood needs to be sufficiently large and diverse to span all relevant concepts. At the same time, the neighborhood needs to be sufficiently local to focus on only those concepts that are relevant to the test point. In Figure 5 (right), we support this hypothesis by varying the neighborhood size for TTT on ImageNet, and finding that the optimal neighborhood size trades off locality and diversity.
Global Local TTT Global TTT
MNIST 98.57 $\pm$ 0.12 99.01 $\pm$ 0.10 36.38 $\pm$ 0.16
ImageNet 78.33 $\pm$ 0.19 79.39 $\pm$ 0.18 77.04 $\pm$ 0.06
Table 2: Accuracy of a linear model trained on the full dataset (global), trained on local neighborhoods (local TTT), and of ten randomly selected local TTT heads evaluated on the entire test set (global TTT). We select ten TTT heads to keep the evaluation computationally tractable. The table reports bootstrap standard errors.
@@
Outlook: Scaling mixtures of experts.
@@ 5 Why may specialization help?
Based on our mechanistic understanding gained in Section 3 through observations O1-O3, we next provide theoretical evidence that the LRH supports our practical observations from Section 4, therefore, potentially offering an understanding for why specialization is effective. To simplify the analysis, we consider a univariate regression setting with input space $\mathcal{X}$ and output space $\mathbb{R}$. Based on the LRH, we posit the existence of an $s$-sparse high-dimensional concept space $\Phi:\mathcal{X}\to\mathbb{R}^{d_{1}}$, and that the ground truth output is a linear combination of concepts, i.e., $\langle\Phi(x),w_{\star}\rangle$ for an arbitrary weight vector $w_{\star}\in\mathbb{R}^{d_{1}}$. Let us denote the learned feature map by $\Psi:\mathcal{X}\to\mathbb{R}^{d_{2}}$ and assume that the feature map is “underparameterized”, i.e., $1\leq d_{2}\ll d_{1}$. We assume access to a training set $\mathcal{D}$ of size $N$.
@@ 5.1 Analyzing the in-distribution test error of TTT
Informally, the LRH states that the model represents high-level concepts linearly as directions in some concept space, with the unembedding $w_{\star}$ defining the “meaning” of concepts. As is well-known from compressed sensing (Candes & Tao, 2006; Donoho, 2006), the feature map $\Psi$ can represent exponentially many concepts in superposition (Elhage et al., 2022), i.e., $d_{1}\sim s\exp(d_{2}/s)$. However, as we will see in this section, even if the underparameterized feature map encodes the structure of the concept space, it is not straightforward to learn a linear mapping of these superimposed concepts. In contrast, we show in an idealized setting based on our empirical observations, that TTT can efficiently learn the meaning of exponentially many concepts from data, by specializing the model to the concepts relevant to the test data. We begin by restating our three key hypotheses based on observations from Section 3. We formalize these hypotheses in Appendix B and include proofs in Appendix C.
@@
In Section 3, we have made the surprising observation that the neighborhood of a test point $x^{\star}$ is explained by only a few active concepts (cf. O2). Based on this, we assume that locally there exists an $\Theta(s)$-sparse concept vector $w_{x^{*}}\in\mathbb{R}^{d_{1}}$ that approximates $\langle\Phi(x),w_{\star}\rangle$ over the test point’s neighborhood. This assumption may seem unrealistic at first, since the neighborhood of $x^{\star}$ has $k$ samples and up to $s\cdot k$ active concepts. Yet, as we observe empirically in Section 3, TTT learns a $\Theta(s)$-sparse regressor that does well on the entire neighborhood.
@@
We expand on our theoretical results in the appendix as follows: In addition to increasing the effective model capacity, we also explore whether one can understand TTT under the LRH through the lens of statistical learning theory. Specifically, in Section A.1, we explore this direction using notions from low-degree polynomials and hypercontractivity (Klivans et al., 2008; Paouris et al., 2022; Damian et al., 2024; Bizeul & Klartag, 2025, and references therein). Moreover, in Section A.2, we contrast TTT to classical non-parametric methods (Fix & Hodges Jr., 1951; Nadaraya, 1964; Watson, 1964) such as majority voting, which underperform in our experiments (cf. Section 4).

### v2

@@ 1 Introduction
We propose to model this phenomenon under the linear representation hypothesis (LRH; Mikolov et al., 2013; Park et al., 2024; 2025), which postulates that models represent high-level concepts—meaningful semantic features—as directions in a latent space. This latent concept space is typically assumed to be sparse conditioned on the input, meaning that any given input activates only a few possible concepts. For a particular prediction task, the linear coefficients of the active concepts effectively characterize their “meaning” for that task. Because the number of real-world concepts far exceeds a model’s dimensionality, these concepts are superimposed within the model’s dense activations (cf. Figure˜1; Elhage et al., 2022). The LRH has been used extensively in prior work on interpretability (Kim et al., 2018) and activation steering (Bolukbasi et al., 2016; Templeton et al., 2024) of foundation models. In this work, we analyze a model where TTT can learn the meaning of these superimposed concepts from data more efficiently than training a “global” model or non-parametric methods.
@@
The LRH can be formalized as follows. We assume the existence of an $s$-sparse concept space $\Phi:\mathcal{X}\to\mathbb{R}^{d_{1}}$} which is approximated by a learned feature map $\Psi:\mathcal{X}\to\mathbb{R}^{d_{2}}$} with $d_{2}\ll d_{1}$. The ground truth to be learned is linear in the concept space, i.e., $f(x)=\langle\Phi(x),w_{\star}\rangle$ for some unknown $w_{\star}\in\mathbb{R}^{d_{1}}$. [^1] In classification, the logits are canonically parameterized as $\langle\Phi(x),w_{c}\rangle$ where $w_{c}$ is the class-specific weight vector. The class probability is then given by $\mathbb{P}(c\mid x)\propto\exp(\langle\Phi(x),w_{c}\rangle)$. In Section˜3, we leverage the LRH to develop a mechanistic understanding of how TTT behaves, and make the following key observations:
@@ 3. O3:
Based on the LRH and our key observations, in Sections˜4 and 5, we analyze when and why TTT can be effective. We find that TTT improves predictions in the underparameterized regime, but its improvement diminishes as models become overparameterized. We support this finding empirically and theoretically:
@@ 2 Related work
So far it has not been well understood why and when TTT is effective. While many different methods have been proposed for TTT, we focus here on analyzing “semi-parametric” TTT (e.g., Hardt & Sun, 2024; Hübotter et al., 2025), where a pre-trained model is fine-tuned with a supervised loss on a small neighborhood of the test point in the training data. This is different from some other methods for test-time “adaptation”, which are commonly applied with distribution shifts (e.g., Wang et al., 2021; Zhang et al., 2022; Durasov et al., 2025). Basu et al. (2023) consider a similar setting to ours, but analyze it through the lens of non-parametric estimation, relying on the smoothness of the target function in the feature space $\Psi$. In contrast, our framework explicitly models the underlying sparse concept space $\Phi$. This explains why TTT substantially outperforms “non-parametric" methods even when the function is locally high-dimensional ($s$-sparse) in the concept space. Furthermore, while most prior theoretical work simply assumes the TTT gradient aligns with the gradient on the oracle label (e.g., Sun et al., 2020), our work provides an idealized model where this alignment is justified.
@@ 3 How does specialization behave?
We begin in this section by developing a mechanistic understanding of how TTT behaves. Since we do not have access to the “true” hypothesized concept space $\Phi$, we train SAEs to learn a concept space $\smash{\hat{\Phi}}$ whose characteristics can be probed. We employ a top-$k$ SAE (Gao et al., 2025) to learn sparse feature representations. After a brief description of the experimental setup, we detail our key observations O1-O3 from Section˜1.
@@ 3.1 Experimental setup
We use the ImageNet-1K dataset (Deng et al., 2009). The dense vectors $\Psi(x)$ are normalized CLIP embeddings (Radford et al., 2021) of dimension $d_{2}=512$ (we use only the <CLS> component). We set the sparse dimension to $d_{1}=8\times d_{2}=4096$ and the sparsity level to $s=16$. For our analysis, we use the SAE’s reconstructions $\smash{\hat{\Psi}(x)}$ rather than the original CLIP embeddings $\Psi(x)$. This choice aligns our experiments more closely with our theoretical model and circumvents known challenges in training SAEs on raw, complex embeddings. This comes at the cost of a mild $6\%$ drop in accuracy for a global linear classifier trained on the embeddings. [^3] Note that the SAE is trained in an unsupervised way, without explicitly retaining classification accuracy, yet using the SAE’s features leads only to a minor drop in accuracy. In Section˜E.2, we additionally validate the SAE framework on MNIST (LeCun et al., 1998b), where we use $\Psi(x)$ directly rather than its reconstruction, as recovering an accurate $\smash{\hat{\Phi}(x)}$ is much easier on this simpler dataset. We include training details in Section˜E.3.
@@ 3.2 Results
Our first hypothesis is that the SAE mapping preserves the angular relationships between a point and its neighbors. To test this, we select a neighborhood for a test point $x^{\star}$ in three different spaces: the original CLIP space ($\Psi$), the reconstructed space ($\smash{\hat{\Psi}}$), and the estimated concept space ($\smash{\hat{\Phi}}$). We then measure the average cosine similarity in the estimated concept space between $x^{\star}$ and points in each neighborhood. As shown in Figure˜2, the distributions of cosine similarities are nearly identical regardless of the space used for neighbor selection. This suggests that the SAE projection to the concept space preserves the local geometric structure, supporting our first key observation.
O2: Neighborhoods are supported by few concepts.
@@
With a sparsity penalty of $\lambda=0.2$, the learned masks are highly sparse, activating on average only $\|m\|_{0}\approx 40$ concepts. This is substantially smaller than the total number of unique concepts active across the neighborhood, which is approximately $180$. As shown in Table˜1 (TTT column), this sparsely supported model performs on par with TTT on top of dense reconstructions $\smash{\hat{\Psi}(x)}$ [^5] TTT in the concept space without extra masking achieves similar results. The fact that it also works with masking provides strong evidence that TTT adjusts only a few selected concepts in the neighborhood.. This suggests that a small, adaptively chosen set of concepts is sufficient to capture the relevant information within a local region. We obtain similar results for the Gemma Scope SAE (Lieberum et al., 2024) on MNIST data, which we present in Section˜E.2.
@@
Figure 3: Comparison of predicted class probabilities for TTT models trained on dense reconstructions ($\smash{\hat{\Psi}}$) and sparse concepts ($\smash{\hat{\Phi}}$), relative to their magnitude. The small relative TV distance ($\ll\!\!1$; defined in Section˜E.1) between both distributions indicates strong functional agreement between TTT in $\smash{\hat{\Psi}}$ and $\smash{\hat{\Phi}}$ space. We show 90% bootstrap confidence intervals across $1000$ test points.
@@
While the adaptive masking in Equation˜3 explicitly enforces a sparse solution, we find evidence that standard TTT in the feature space implicitly favors a solution that is sparse in the concept space. First, the TTT models trained on dense reconstructions $\smash{\hat{\Psi}(x)}$ and sparse concepts $\smash{\hat{\Phi}_{m}(x)}$ achieve nearly identical accuracy (cf. Table 1). Furthermore, their predictions agree in $\approx\!\!89\%$ of cases, indicating that they learn functionally equivalent classifiers (apart from pathological examples). Figure˜3 reinforces this by showing that both models lead to closely matched predictive distributions over the top-10 predicted classes. In Figure˜3, we compare the ordered predicted probabilities for $\smash{\hat{\Phi}}$ to the corresponding probabilities for $\smash{\hat{\Psi}}$, matching the distributions’ temperatures, and averaging over all test points. Their strong correspondence suggests that TTT on reconstructed embeddings is implicitly biased towards a sparse solution in the underlying concept space. This phenomenon may be linked to the implicit bias of optimization algorithms (e.g., SGD or Adam), which are known to favor minimum-norm solutions (Gunasekar et al., 2018; belkin2019reconciling; frei2022benign). When the feature map superimposes concepts, this implicit bias may favor sparse solutions in the underlying concept space (Vaskevicius et al., 2019).
@@ 4 When does specialization help?
After gaining some mechanistic understanding of TTT in Section˜3, we next study when specialization through TTT improves over a globally trained model. In Section˜5, we then relate our results from Sections 3-4, by providing theoretical support for why the LRH and our observations O1-O3 may support our practically relevant findings from this section.
@@ 4.1 Settings
We consider the following three tasks, and provide implementation details in Appendix˜E:
  • MNIST. We train a global classifier following the LeNet-5 (LeCun et al., 1998a) architecture with a cross-entropy loss on the MNIST (LeCun et al., 1998b) dataset of handwritten digits. To distinguish learning of concepts from TTT’s ability to identify the correct weights for each locally relevant concept, we train only the last linear layer during TTT. The last linear layer is updated to minimize the cross-entropy loss on the neighborhood of a test point (cf. Equation˜2). We report classification error on the test set.
@@ 4.2 Results
We perform a scaling study by varying the model size and comparing the performance of TTT against global training and a majority vote on the neighborhood (i.e., a simple non-parametric baseline). Figure˜4 shows the results. In MNIST, we train convolutional neural networks of varying sizes as summarized in Appendix˜E. In ImageNet, we train multi-layer perceptrons on top of CLIP embeddings, varying the hidden dimension. In language modeling, we evaluate Qwen2.5 base models of sizes ranging from 0.5B to 32B parameters. We find across all tasks that TTT outperforms global training and majority vote, with the performance gap shrinking as the model size increases. We hypothesize that at a larger model size, fewer concepts have to be superimposed in latent space, leading to less interference when globally mapping latent representations to predictions. While a larger model size allows for better global disentanglement of concepts, TTT can compensate for limited model capacity by adapting the head to the specific concepts in a local neighborhood.
@@
Next to performing a scaling study on model size, we also vary the dataset size. Figure˜5 shows the results at a fixed model scale. We subsample the training datasets of ImageNet and MNIST to fractions ranging from $1\%$ to $100\%$ of the original training set size, ensuring that all subsampled datasets are class-balanced. We then train global models and evaluate TTT on the respective test sets. We find that TTT consistently outperforms global training, with the slight trend of the performance gap widening as the dataset size increases. We hypothesize that larger datasets provide richer neighborhoods for local adaptation, enabling TTT to specialize more effectively to the specific concepts relevant to each test point.
@@
To validate that TTT improves predictions locally, we globally evaluate some TTT heads. As shown in Table˜2, while improving accuracy on the test point and neighborhood, the global accuracy of these fixed TTT heads is significantly lower than that of the model trained on the entire dataset without TTT. These results confirm that, while TTT can provide localized performance improvements when adapted individually at test time, such benefits do not generalize when the same adaptation is applied globally. We further hypothesize that the neighborhood needs to be sufficiently large and diverse to span all relevant concepts. At the same time, the neighborhood needs to be sufficiently local to focus on only those concepts that are relevant to the test point. In Figure 5 (right), we support this hypothesis by varying the neighborhood size for TTT on ImageNet, and finding that the optimal neighborhood size trades off locality and diversity.
@@
Global TTT on Test Sample TTT on Neighborhood Global TTT
MNIST 98.57 $\pm$ 0.12 99.01 $\pm$ 0.10 100.00 $\pm$ 0.00 36.38 $\pm$ 0.16
ImageNet 78.33 $\pm$ 0.19 79.39 $\pm$ 0.18 95.19 $\pm$ 0.00 77.04 $\pm$ 0.06
Table 2: Accuracy of a linear model trained on the full dataset (global), TTT evaluated on the test sample, TTT evaluated on the neighborhood, and of ten randomly selected local TTT heads evaluated on the entire test set (global TTT). We select ten TTT heads to keep the evaluation computationally tractable. The table reports bootstrap standard errors.
@@ 5 Why may specialization help?
Based on our mechanistic understanding gained in Section˜3 through observations O1-O3, we next provide theoretical evidence that the LRH supports our practical observations from Section˜4, therefore, potentially offering an understanding for why specialization is effective. To simplify the analysis, we consider a univariate regression setting with input space $\mathcal{X}$ and output space $\mathbb{R}$. Based on the LRH, we posit the existence of an $s$-sparse high-dimensional concept space $\Phi:\mathcal{X}\to\mathbb{R}^{d_{1}}$, and that the ground truth output is a linear combination of concepts, i.e., $\langle\Phi(x),w_{\star}\rangle$ for an arbitrary weight vector $w_{\star}\in\mathbb{R}^{d_{1}}$. Let us denote the learned feature map by $\Psi:\mathcal{X}\to\mathbb{R}^{d_{2}}$ and assume that the feature map is “underparameterized”, i.e., $1\leq d_{2}\ll d_{1}$. We assume access to a training set $\mathcal{D}$ of size $N$.
@@ 5.1 Analyzing the in-distribution test error of TTT
Informally, the LRH states that the model represents high-level concepts linearly as directions in some concept space, with the unembedding $w_{\star}$ defining the “meaning” of concepts. As is well-known from compressed sensing (Candes & Tao, 2006; Donoho, 2006), the feature map $\Psi$ can represent exponentially many concepts in superposition (Elhage et al., 2022), i.e., $d_{1}\sim s\exp(d_{2}/s)$. However, as we will see in this section, even if the underparameterized feature map encodes the structure of the concept space, it is not straightforward to learn a linear mapping of these superimposed concepts. In contrast, we show in an idealized setting based on our empirical observations, that TTT can efficiently learn the meaning of exponentially many concepts from data, by specializing the model to the concepts relevant to the test data. We begin by restating our three key hypotheses based on observations from Section˜3. We formalize these hypotheses in Appendix˜B and include proofs in Appendix˜C.
@@
In Section˜3, we have made the surprising observation that the neighborhood of a test point $x^{\star}$ is explained by only a few active concepts (cf. O2). Based on this, we assume that locally there exists an $\Theta(s)$-sparse concept vector $w_{x^{*}}\in\mathbb{R}^{d_{1}}$ that approximates $\langle\Phi(x),w_{\star}\rangle$ over the test point’s neighborhood. This assumption may seem unrealistic at first, since the neighborhood of $x^{\star}$ has $k$ samples and up to $s\cdot k$ active concepts. Yet, as we observe empirically in Section˜3, TTT learns a $\Theta(s)$-sparse regressor that does well on the entire neighborhood.
@@
We expand on our theoretical results in the appendix as follows: In addition to increasing the effective model capacity, we also explore whether one can understand TTT under the LRH through the lens of statistical learning theory. Specifically, in Section˜A.1, we explore this direction using notions from low-degree polynomials and hypercontractivity (Klivans et al., 2008; Paouris et al., 2022; Damian et al., 2024; Bizeul & Klartag, 2025, and references therein). Moreover, in Section˜A.2, we contrast TTT to classical non-parametric methods (Fix & Hodges Jr., 1951; Nadaraya, 1964; Watson, 1964) such as majority voting, which underperform in our experiments (cf. Section˜4).